# DReSS: Data-driven Regularized Structured Streamlining for Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have achieved significant progress across various domains, but their increasing scale results in high computational and memory costs. Recent studies have revealed that LLMs exhibit sparsity, providing the potential to reduce model size through pruning techniques. However, existing pruning methods typically follow a prune-then-finetune paradigm. Since the pruned components still contain valuable information, their direct removal often leads to irreversible performance degradation, imposing a substantial computational burden to recover performance during finetuning. In this paper, we propose a novel paradigm that first applies regularization, then prunes, and finally finetunes. Based on this paradigm, we introduce DReSS, a simple and effective Data-driven Regularized Structured Streamlining method for LLMs. By leveraging a small amount of data to regularize the components to be pruned, DReSS explicitly transfers the important information to the remaining parts of the model in advance. Compared to direct pruning, this can reduce the information loss caused by parameter removal, thereby enhancing its language modeling capabilities. Experimental results demonstrate that DReSS significantly outperforms existing pruning methods even under extreme pruning ratios, significantly reducing latency and increasing throughput.
@@ 1 Introduction
Large language models (LLMs) have achieved significant advancements across a wide range of tasks and domains, demonstrating their robust capabilities (Zhang et al., 2022; Achiam et al., 2023; Touvron et al., 2023; Wu et al., 2024). However as the model size grows, the increased number of parameters results in substantial computational and memory demands. These factors severely hinder the deployment of LLMs in practical applications. Therefore, it is urgent to develop methods that can reduce model size while maintaining performance.
@@
To address these challenges, several methods have been proposed, including pruning (Frantar & Alistarh, 2023; Sun et al., 2023; Ma et al., 2023; An et al., 2024), quantization (Frantar et al., 2022; Xiao et al., 2023), knowledge distillation (Shridhar et al., 2022; Hsieh et al., 2023), and low-rank decomposition (Saha et al., 2023). In this work, we mainly focus on pruning, an efficient and highly generalizable method that can be combined with other model compression strategies. Pruning techniques are generally classified into two primary categories: unstructured pruning (Frantar & Alistarh, 2023; Sun et al., 2023) and structured pruning (Ma et al., 2023; An et al., 2024). Compared to unstructured pruning, structured pruning offers the flexibility to do recovery fine-tuning (RFT) on the pruned model for specific downstream tasks without relying on specialized hardware (Zhu et al., 2024). Moreover, the model obtained through structured pruning typically achieves much faster inference speed, while ensuring the same perplexity and accuracy as the model obtained through unstructured pruning.
@@
Despite these advancements, existing structured pruning methods still have some limitations. They all follow the paradigm of first selecting channels or layers for pruning based on a designed metric, and then performing RFT (Chavan et al., 2024). However, they neglect the fact that important information can also exist in the pruned channels or layers (Dettmers et al., 2022; Xiao et al., 2023; Yin et al., 2023), and directly removing them leads to an irreversible decline in model performance. Thus, in order to achieve satisfactory performance with the pruned model, a large amount of data must be introduced for RFT (Ma et al., 2023), which increases computational overhead. Moreover, they often lead to a collapse in model performance at high pruning ratios, limiting their effectiveness in reducing lantency and enhancing throughput.
@@
To tackle these limitations, in this paper, we propose a new pruning paradigm that first applies regularization to the components to be pruned, then removes those parameters, and finally performs RFT. To the best of our knowledge, we are the first to propose this paradigm. The regularization process explicitly transfers important information from the parts to be pruned to the remaining parts, thereby reducing the information loss caused by direct parameter removal. Based on this paradigm, we introduce DReSS, a Data-driven Regularized Structured Streamlining method for LLMs. The comparison of the proposed method with existing channel-wise pruning approaches is shown in Figure 1. Specifically, the process of DReSS is divided into the following steps: (1) select data for pre-pruning regularization and post-pruning RFT, (2) apply regularization to the selected channels in the parameter matrix, and (3) prune the specified channels, then (4) do RFT. First, we randomly selecting a small subset of data from widely used benchmark datasets, as the amount of selected data is small, the overhead in the regularization and RFT process can be significantly reduced. Second, we determine the channels to be pruned based on the pruning ratio, and apply regularization to these channels using the selected data. This shifts important information to the unpruned parts, significantly mitigating performance degradation caused by direct pruning and ensuring the model maintains strong performance even at high pruning ratios, thus enabling substantial inference speedup. Third, we prune the parameters that have been regularized. Finally, we do RFT on the pruned model using a small subset of samples selected in the first step. Extensive experiments demonstrate that, DReSS significantly surpasses existing pruning methods in terms of perplexity and accuracy across various pruning ratios, it also substantially improves inference speed. The main contributions of our work are summarized as follows:
@@
  • Propose A New Paradigm: By first applying regularization, then pruning, and finally performing RFT, DReSS reduces information loss from direct parameter removal, enhancing the model’s language modeling ability and lowering the computational cost of RFT.
@@
  • High Performance: DReSS outperforms existing structured pruning methods on both generation and zero-shot tasks. The pruned model shows a significant increase in throughput and a notable reduction in latency compared to the dense model.
@@
  • Low Overhead: By using only a small amount of data during the regularization and optional RFT process, the pruning overhead is minimal.
@@
Figure 1: A comparison between existing channel-wise structured pruning methods and DReSS. The deeper blue square represents greater performance impact; the red exclamation mark indicates the important information. The taller orange-yellow cylinder represents larger data volume. Here, L($\cdot$) represents the regularization, and S represents pruning. Above: Existing methods first select channels based on importance, followed by pruning and RFT on the pruned model. Below: DReSS first regularizes channels to minimize performance impact, prunes the channels, then do RFT with a little amount of data.
@@ 2.1 Pruning Methods
Unstructured pruning methods remove individual weights based on their importance. For instance, Magnitude pruning (Han et al., 2015; Zhou et al., 2021) assumes larger weights are more important, Wanda (Sun et al., 2023) uses the product of layer weights and activations as the metric, and SparseGPT (Frantar & Alistarh, 2023) relies on the weight squared divided by the inverse Hessian matrix. The main advantage is flexibility, enabling the removal of unimportant weights while retaining critical ones, allowing high sparsity in LLMs. However, its limitations include the need for specialized hardware for accelerated inference (Xia et al., 2023a), high sparsity for significant acceleration (Wang, 2020), and inability to perform RFT on downstream tasks.
@@ 2.2 Regularization
Regularization is widely employed in machine learning (Hoerl & Kennard, 1970; Poggio et al., 1987; Balestriero et al., 2022), such as feature selection (Tibshirani, 1996) and preventing model overfitting (Santos & Papa, 2022). The $\ell_{1}$-norm typically induces sparsity by driving certain coefficients to zero, while $\ell_{2}$-norm encourages smoother solutions (Boyd & Vandenberghe, 2004). Both of them can significantly alter the distribution pattern of the data (Han et al., 2015; Tao et al., 2023). Motivated by this observation, we propose a data-driven regularization approach for parameters designated for pruning. By applying a small amount of data, the regularization probably transfers important information from the pruned parameter space to the retained parts of the model, thereby enhancing the performance of the pruned model.
@@ 3 Methodology
  • Regularization: Applying regularization to the selected channels of weight matrices during both forward and backward propagations.
@@
     update $W$ using BP algorithm
@@ 3.1 Data Selection
The required data is selected from widely used calibration datasets, including Alpaca (Taori et al., 2023), WikiText-2 (Merity et al., 2016), PTB (Marcus et al., 1993), and C4 (Raffel et al., 2020). For instance, approximately 1,000 samples are randomly selected from the WikiText-2 training dataset for both pre-pruning regularization and post-pruning RFT. In Section 4.5, we analyze the impact of different datasets.
@@ 3.2 Regularization
To provide a clearer understanding of this section, we first introduce the relevant notations. Let $p$ denote the pruning ratio, which indicates that $p\%$ of the model’s parameters need to be pruned. The model dimension of the LLMs is represented by $d$, and the number of layers is denoted by $l$. Additionally, let $W_{emb}\in\mathbb{R}^{b\times n\times d}$ represents the data feature map, and $W_{pos}\in\mathbb{R}^{b\times n\times d}$ denotes the positional feature map, where $b$ is the batch size and $n$ is the number of tokens. The model weights are represented by $W\in\mathbb{R}^{d_{1}\times d_{2}}$. Specifically, $W_{Q}^{i}\in\mathbb{R}^{d\times d_{1}}$, $W_{K}^{i}\in\mathbb{R}^{d\times d_{1}}$, $W_{V}^{i}\in\mathbb{R}^{d\times d_{1}}$, and $W_{o}^{i}\in\mathbb{R}^{d_{1}\times d}$ represent the parameter matrices in the $i$th layer Attention block. Similarly, $W_{up}^{i}\in\mathbb{R}^{d\times d_{2}}$ and $W_{down}^{i}\in\mathbb{R}^{d_{2}\times d}$ represent the parameter matrices in the $i$th layer FFN block. Additionally, $w_{N}^{i}\in\mathbb{R}^{d}$ denotes the LayerNorm vector in the $i$-th layer, and $W_{lm}\in\mathbb{R}^{b\times n}$ represents the language modeling matrix. Finally, $I\in\mathbb{R}^{d\times d}$ represents the identity matrix, let $R\in\mathbb{R}^{d\times d}$ represent a pseudo-index selection matrix, which is a diagonal matrix consisting of $(1-p\%)\times d$ zeros and $p\%\times d$ ones, where the ones correspond to the selected indices.
@@
(Dependency between the Attention Block and FFN Block.) If regularization is applied to certain columns of $W_{down}^{i-1}$ in the $(i-1)$th layer FFN block, then regularization must also be applied to the corresponding rows of $W_{Q}^{i}$, $W_{K}^{i}$, and $W_{V}^{i}$ in the $i$th layer Attention block. The same dependency applies to $W_{o}^{i}$ and $W_{up}^{i-1}$.
@@
(If the loss function includes the $\ell_{1}$-norm, it can also be solved using backpropagation.) The following unconstrained optimization problem is equivalent to the constrained optimization problem, where $|\cdot|_{1}$ denotes the $\ell_{1}$-norm.
$\min\quad|x|_{1}\quad\iff\quad\min_{x,y}$ $\quad\mathbf{1}^{T}y$ (5)
@@
We formalize the minimization of the overall loss as an optimization problem. When $\ell_{2}$-norm is used, the problem can be efficiently solved using backpropagation. In contrast, when the $\ell_{1}$-norm is employed, as the regularization losses introduced by various components in LLMs are similar, we consider only ~$\tilde{\mathcal{L}}_{\text{FFN}}^{i}$ as an example. According to Proposition 3.2, the optimization problem in Equation 6 can be equivalently transformed into a constrained formulation in Equation 7, enabling its solution by backpropagation.
$\min_{W}\quad\mathcal{L}(W,X)+\lambda|RW_{up}^{i}|_{1}+\lambda|W_{down}^{i}R|_{1}$ (6)
@@
During the forward and backward propagation processes, due to the application of regularization on certain specific channels in LLMs, the values within these rows or columns are significantly reduced. Intuitively, this reduction could potentially decrease the impact of these parameters on the model’s performance.
@@ 3.4 Optional RFT
After pruning, we perform RFT on the model using a small subset of the data selected in Section 3.1, leveraging LoRA (Hu et al., 2021). This step is optional, as experimental results in section 4.7 show that DReSS without RFT still maintains competitive performance. in terms of both perplexity and zero-shot accuracy.
Table 1: Performance comparison of different pruning methods. ’PPL’ refers to the perplexity on WikiText-2. The accuracy is reported on five zero-shot tasks: PIQA, WinoGrande, HellaSwag, ARC-e, and ARC-c, as well as the average accuracy. The best result is highlighted in bold, and the second-best in underlined. DReSS-$\ell_{2}$ denotes the use of the $\ell_{2}$-norm, while DReSS-$\ell_{1}$ denotes the use of the $\ell_{1}$-norm.
@@ 4 Experiments
This section introduce experimental setup 4.1 and analyzes the effectiveness of DReSS from the following aspects: performance 4.2, acceleration effectiveness 4.3, robustness under different pruning ratios 4.4, robustness under different datasets 4.5, low overhead 4.6, and ablation study 4.7.
@@ 4.1 Experimental Setup
Implementation DReSS and baseline methods are implemented on the PyTorch framework (Paszke et al., 2019), utilizing the Hugging Face Transformers library (Wolf, 2019). All experiments are conducted on 80GB NVIDIA A100 GPUs. For fairness, we use llm-eval-harness (Gao et al., 2024a) to evaluate the pruned models.
Datasets To evaluate the performance of DReSS, we conducted experiments on generation and zero-shot tasks. For generation task, following prior work(Ashkboos et al., 2024), we evaluate the model’s perplexity on WikiText-2. For zero-shot task evaluation, the benchmarks consists of PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), HellaSwag (Zellers et al., 2019), ARC-e and ARC-c (Clark et al., 2018). To demonstrate that DReSS does not have a specific dependency on the regularization data, we use a random selection of data from Alpaca (Taori et al., 2023), WikiText-2 (Merity et al., 2016), PTB (Marcus et al., 1993), and C4 (Raffel et al., 2020) for calibration prior to pruning and for the RFT after pruning.
Models The models pruned using DReSS include the LLaMA models (LLaMA2-7B, LLaMA2-13B) (Touvron et al., 2023), OPT models (OPT-2.7B, OPT-6.7B, OPT-13B) (Zhang et al., 2022), and Phi-2 (Javaheripi et al., 2023).
Baselines We evaluate DReSS against structured pruning methods: LLM Surgeon (van der Ouderaa et al., 2023), SliceGPT (Ashkboos et al., 2024), and SLEB (Song et al., 2024). Among them, LLM Surgeon and SliceGPT are channel-wise, SLEB is layer-wise. To ensure a fair comparison, in the subsequent experiments, we ensure that the data used for regularization in DReSS is consistent with the calibration data used by LLM Surgeon, SliceGPT, and SLEB before pruning, and that the data used for RFT in each method is also the same.
Evaluation Metrics The performance on generation task is measured by perplexity, a well-established and robust metric (Yao et al., 2022), while zero-shot tasks performance is evaluated using accuracy (Dong et al., 2024). The acceleration effects are represented by throughput and latency (Song et al., 2024).

### v2

@@ Abstract
Large language models (LLMs) have achieved significant progress across various domains, but their increasing scale leads to high computational and memory costs. Recent studies show that LLMs exhibit sparsity, which can be exploited for pruning. However, existing pruning methods typically follow a prune-then-finetune paradigm. Since the pruned components still contain valuable information, their direct removal often leads to irreversible performance degradation, necessitating expensive fine-tuning to recover performance. To address this, we propose a new paradigm: first apply regularization, then prune, and finally fine-tune. Based on this paradigm, we introduce DReSS, a simple and effective Data-driven Regularized Structured Streamlining method for LLMs. By leveraging a small amount of data to regularize the components before pruning, DReSS transfers the important information to the remaining parts of the model in advance. Compared to direct pruning, this can reduce the information loss caused by parameter removal, thereby enhancing its language modeling capabilities. Experimental results demonstrate DReSS achieves comparable performance to conventional methods even without fine-tuning, drastically alleviating computational costs. Moreover, DReSS significantly outperforms existing pruning methods even under extreme pruning ratios, significantly reducing latency and increasing throughput.
@@ 1 Introduction
Large language models (LLMs) have achieved significant advancements across a wide range of tasks and domains, demonstrating their robust capabilities (Zhang et al., 2022; Achiam et al., 2023; Touvron et al., 2023; Wu et al., 2024). However, as the model size increases, the growing number of parameters leads to significant computational and memory requirements, which significantly hinder the practical deployment of LLMs. Consequently, it is urgent to develop methods that can reduce model size while maintaining performance.
@@
To address these challenges, several methods have been proposed, including pruning (Frantar & Alistarh, 2023; Sun et al., 2023; Ma et al., 2023; An et al., 2024), quantization (Frantar et al., 2022; Xiao et al., 2023), knowledge distillation (Shridhar et al., 2022; Hsieh et al., 2023), and low-rank decomposition (Saha et al., 2023). In this work, we mainly focus on pruning—an efficient and highly generalizable approach that can be seamlessly integrated with other model compression strategies. Pruning techniques are generally classified into two primary categories: unstructured pruning (Frantar & Alistarh, 2023; Sun et al., 2023) and structured pruning (Ma et al., 2023; An et al., 2024). Compared to unstructured pruning, structured pruning offers the flexibility to do recovery fine-tuning (RFT) for specific downstream tasks without relying on specialized hardware (Zhu et al., 2024). Moreover, the model obtained through structured pruning typically achieves much faster inference speed due to the regular data patterns.
@@
Despite these advancements, existing structured pruning methods still have some limitations. They all follow the paradigm of first selecting channels or layers to prune based on a designed metric, and then performing RFT (Chavan et al., 2024). However, they neglect the fact that important information can also exist in the pruned parts (Dettmers et al., 2022; Xiao et al., 2023; Yin et al., 2023), and directly removing them leads to an irreversible decline in model performance. Thus, achieving satisfactory performance with pruned models often requires extensive data for RFT (Ma et al., 2023), increasing computational costs. Additionally, high pruning ratios frequently cause performance collapse, limiting their effectiveness in reducing latency and improving throughput.
@@
To tackle these limitations, in this paper, we propose a new pruning paradigm that first applies regularization to the components to be pruned, then removes those parameters, and finally performs RFT. To the best of our knowledge, we are the first to propose this paradigm. The regularization process explicitly transfers important information from the parts to be pruned to the remaining parts, thereby reducing the information loss caused by direct parameter removal. Based on this paradigm, we introduce DReSS, a Data-driven Regularized Structured Streamlining method for LLMs. The comparison of the proposed method with existing channel-wise pruning approaches is shown in Figure 1. Specifically, the process of DReSS is divided into the following steps: (1) select data for pre-pruning regularization and post-pruning RFT, (2) apply regularization to the selected channels in the parameter matrix, and (3) prune the specified channels, finally (4) do RFT. First, we randomly select a small subset of data from widely used benchmark datasets. Since the amount of selected data is small, the overhead in the regularization and RFT process can be significantly reduced. Second, we determine the channels to be pruned based on the pruning ratio, and apply regularization to these channels using the selected data. This redistributes important information to the unpruned components, thereby significantly mitigating performance degradation caused by direct pruning. As a result, the model can maintain robust performance even at high pruning ratios, facilitating a substantial speedup in inference. Third, we prune the parameters that have been regularized. Finally, we perform RFT on the pruned model using a small subset of samples selected in the first step. Extensive experiments show that DReSS significantly outperforms existing pruning methods in perplexity and accuracy across various pruning ratios, even without RFT, while also greatly enhancing inference speed. The main contributions of our work are summarized as follows:
@@
  • Propose A New Paradigm: By sequentially applying regularization, pruning, and RFT, DReSS minimizes information loss caused by direct parameter removal, thereby improving the model’s language modeling capabilities and reducing the computational cost of RFT.
@@
  • High Performance: DReSS surpasses existing structured pruning methods in both generation and zero-shot tasks. The pruned model achieves a substantial increase in throughput and a significant reduction in latency compared to the dense model.
@@
  • Low Overhead: By leveraging only a small amount of data for regularization and optional RFT, DReSS incurs minimal overhead.
@@
Figure 1: A comparison between existing channel-wise structured pruning methods and DReSS. The deeper blue square represents greater performance impact; the red exclamation mark indicates the important information. The taller orange-yellow cylinder represents larger data volume. L($\cdot$) represents the regularization, and S represents pruning. Above: Existing methods first select channels based on importance, followed by pruning and RFT on the pruned model. Below: DReSS first regularizes channels to transfer important information, prunes the channels, then do RFT with a little amount of data.
@@ 2.1 Pruning Methods
Unstructured pruning methods remove individual weights based on their importance. For instance, Magnitude (Han et al., 2015; Zhou et al., 2021) assumes larger weights are more important, Wanda (Sun et al., 2023) uses the product of layer weights and activations as the metric, and SparseGPT (Frantar & Alistarh, 2023) relies on the weight squared divided by the inverse Hessian matrix. The main advantage is flexibility, enabling the removal of unimportant weights while retaining critical ones, allowing high sparsity in LLMs. However, its limitations include the requirement for specialized hardware to accelerate inference (Xia et al., 2023a), the necessity for high sparsity to achieve substantial acceleration (Wang, 2020), and the inability to perform RFT on downstream tasks.
@@ 2.2 Regularization
Regularization is widely employed in machine learning (Hoerl & Kennard, 1970; Poggio et al., 1987; Balestriero et al., 2022), such as feature selection (Tibshirani, 1996) and preventing model overfitting (Santos & Papa, 2022). The $\ell_{1}$-norm typically induces sparsity by driving certain coefficients to zero, while $\ell_{2}$-norm encourages smoother solutions (Boyd & Vandenberghe, 2004). Both of them can significantly alter the distribution pattern of the data (Han et al., 2015; Tao et al., 2023). Motivated by this observation, we propose a data-driven regularization approach. By applying a small amount of data, the regularization transfers important information from the pruned parameter space to the retained parts of the model, thereby enhancing the performance of the pruned model.
@@ 3 Methodology
  • Regularization: Applying regularization to the selected channels of weight matrices during forward and backward propagations.
@@
     update $W$ using backpropagation algorithm
@@ 3.1 Data Selection
The required data is selected from widely used calibration datasets, including Alpaca (Taori et al., 2023), WikiText-2 (Merity et al., 2016), PTB (Marcus et al., 1993), and C4 (Raffel et al., 2020). For instance, approximately 1,000 samples are randomly selected from the WikiText-2 training dataset for pre-pruning regularization and post-pruning RFT. In Section 4.5, we analyze the impact of using different datasets.
@@ 3.2 Regularization
To provide a clearer understanding of this section, we first introduce the relevant notations. Let $p$ denote the pruning ratio, which indicates that $p\%$ of the model’s parameters need to be pruned. The model dimension of the LLMs is represented by $d$, and the number of layers is denoted by $l$. Additionally, let $W_{emb}\in\mathbb{R}^{b\times n\times d}$ represents the data feature map, and $W_{pos}\in\mathbb{R}^{b\times n\times d}$ denotes the positional feature map, where $b$ is the batch size and $n$ is the number of tokens. The model weights are represented by $W\in\mathbb{R}^{d_{1}\times d_{2}}$. Specifically, $W_{Q}^{i}\in\mathbb{R}^{d\times d_{1}}$, $W_{K}^{i}\in\mathbb{R}^{d\times d_{1}}$, $W_{V}^{i}\in\mathbb{R}^{d\times d_{1}}$, and $W_{o}^{i}\in\mathbb{R}^{d_{1}\times d}$ represent the parameter matrices in the $i$th layer Attention block. Similarly, $W_{up}^{i}\in\mathbb{R}^{d\times d_{2}}$ and $W_{down}^{i}\in\mathbb{R}^{d_{2}\times d}$ represent the parameter matrices in the $i$th layer FFN block. Additionally, $w_{N}^{i}\in\mathbb{R}^{d}$ denotes the LayerNorm vector in the $i$th layer, and $W_{lm}\in\mathbb{R}^{d\times b\times n}$ represents the language modeling matrix. Finally, $I\in\mathbb{R}^{d\times d}$ represents the identity matrix, let $R\in\mathbb{R}^{d\times d}$ represent a pseudo-index selection matrix, which is a diagonal matrix consisting of $(1-p\%)\times d$ zeros and $p\%\times d$ ones, where the ones correspond to the selected indices.
@@
(Dependency between the Attention Block and FFN Block. Proof in Appendix A.1). If regularization is applied to certain columns of $W_{down}^{i-1}$ in the $(i-1)$th layer FFN block, then regularization must also be applied to the corresponding rows of $W_{Q}^{i}$, $W_{K}^{i}$, and $W_{V}^{i}$ in the $i$th layer Attention block. The same dependency applies to $W_{o}^{i}$ and $W_{up}^{i-1}$.
@@
(If the loss function includes the $\ell_{1}$-norm, it can also be solved using backpropagation. Proof in Appendix A.2). The following unconstrained optimization problem is equivalent to the constrained optimization problem, where $\|\cdot\|_{1}$ denotes the $\ell_{1}$-norm.
$\min\quad||x||_{1}\quad\iff\quad\min_{x,y}$ $\quad\mathbf{1}^{T}y$ (5)
@@
We formalize the minimization of the overall loss as an optimization problem. When $\ell_{2}$-norm is used, the problem can be efficiently solved using backpropagation (BP) algorithm (Rumelhart et al., 1986). In contrast, when the $\ell_{1}$-norm is employed, as the regularization losses introduced by various components in LLMs are similar, we consider only ~$\tilde{\mathcal{L}}_{\text{FFN}}^{i}$ as an example. According to Proposition 3.2, the optimization problem in Equation 6 can be equivalently transformed into a constrained formulation in Equation 7, enabling its solution by BP algorithm.
$\min_{W}\quad\mathcal{L}(W,X)+\lambda||RW_{up}^{i}||_{1}+\lambda||W_{down}^{i}R||_{1}$ (6)
@@
During the forward and backward propagation processes, due to the application of regularization on certain specific channels in LLMs, the values within these rows or columns are significantly reduced. Intuitively, this reduction could potentially transfer important information and decrease the impact of these parameters on the model’s performance.
@@ 3.4 Optional RFT
After pruning, we perform RFT on the model using a small subset of the data selected in Section 3.1, leveraging LoRA (Hu et al., 2021). This step is optional, as experimental results in section 4.7 show that DReSS without RFT still maintains competitive performance in terms of both perplexity and zero-shot accuracy.
Table 1: Performance comparison of different pruning methods. ‘PPL’ refers to the perplexity on WikiText-2. The accuracy is reported on five zero-shot tasks: PIQA, WinoGrande, HellaSwag, ARC-e, and ARC-c, as well as the average accuracy. The best result is highlighted in bold, and the second-best in underlined. DReSS-$\ell_{2}$ denotes the use of the $\ell_{2}$-norm, while DReSS-$\ell_{1}$ denotes the use of the $\ell_{1}$-norm.
@@ 4 Experiments
This section introduces experimental setup (4.1) and analyzes the effectiveness of DReSS from the following aspects: performance comparison (4.2), acceleration (4.3), robustness under different pruning ratios (4.4), dependency on different datasets (4.5), low overhead (4.6), and ablation study (4.7).
@@ 4.1 Experimental Setup
Implementation: DReSS and baseline methods are implemented on the PyTorch framework (Paszke et al., 2019) , utilizing the Hugging Face Transformers library (Wolf, 2019). All experiments are conducted on 80GB NVIDIA A100 GPUs. For fairness, we use llm-eval-harness (Gao et al., 2024a) to evaluate the pruned models. More details are shown in Appendix B.
Datasets: To evaluate the performance of DReSS, we conducted experiments on generation and zero-shot tasks. For generation task, following prior work(Ashkboos et al., 2024), we evaluate the model’s perplexity on WikiText-2. For zero-shot task evaluation, the benchmarks consists of PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), HellaSwag (Zellers et al., 2019), ARC-e and ARC-c (Clark et al., 2018). To demonstrate that DReSS is not specifically dependent on the regularization data, we use a randomly selected subset of data from Alpaca (Taori et al., 2023), WikiText-2 (Merity et al., 2016), PTB (Marcus et al., 1993), and C4 (Raffel et al., 2020) for calibration before pruning and for RFT after pruning.
Models: The models pruned using DReSS include the LLaMA models (LLaMA2-7B, LLaMA2-13B) (Touvron et al., 2023), OPT models (OPT-2.7B, OPT-6.7B, OPT-13B) (Zhang et al., 2022), and Phi-2 (Javaheripi et al., 2023).
Baselines: We evaluate DReSS against competitive structured pruning methods: LLM Surgeon (van der Ouderaa et al., 2023), SliceGPT (Ashkboos et al., 2024), and SLEB (Song et al., 2024). Among them, LLM Surgeon and SliceGPT are channel-wise, SLEB is layer-wise. To ensure a fair comparison, in the subsequent experiments, we ensure that the data used for regularization in DReSS is consistent with the calibration data used by LLM Surgeon, SliceGPT, and SLEB before pruning, and that the data used for RFT in each method is also the same. We also compared with powerful unstructured pruning methods like Mangniude (Han et al., 2015), Wanda (Sun et al., 2023), and SparseGPT (Frantar & Alistarh, 2023) and the results are in Appendix C.
Evaluation Metrics: The performance on generation task is measured by perplexity, a well-established and robust metric (Yao et al., 2022), while zero-shot tasks performance is evaluated using accuracy (Dong et al., 2024). The acceleration effects are represented by throughput and latency (Song et al., 2024).